x^2-2x=4x+16
for x

Answer by algebrapro18(249)

x^2-2x=4x+16
x^2-2x-4x-16=0
x^2-6x-16=0
(x-8)(x+2)=0
x-8=0 x+2=0
x=8 x= -2